# EFKAN: A KAN-Integrated Neural Operator For Efficient Magnetotelluric Forward Modeling

### v1

@@ 3 Computational Experiments
In this section, we use synthetic data to evaluate the effectiveness of the proposed method. To make the experiments as close as possible to field MT forward modeling, we utilize the Gaussian random field (GRF)-based approach to simulate resistivity models for training and testing the EFKAN. We present a series of computational experiments comparing the efficiency and accuracy of EFNO and EFKAN in solving 2-D MT forward problems. Specifically, we focus on using EFKAN to improve the accuracy of solutions for the 2-D Helmholtz equation and explore its applicability to spatial-temporal coordinates and frequencies not included in the training data, as well as its performance on small-scale datasets. All experiments are implemented with the PyTorch platform, and we train and test EFKAN and its competitive approach on a NVIDIA Tesla K80 GPU. It should be pointed out that the colorbars associated with EFKAN errors differ from those of EFNO errors in this study. Consequently, when comparing the prediction errors of EFNO and EFKAN, it is essential to take into account the differences between their respective colorbars.
@@ 3.1 Data Generation
To make the conductivity $\sigma$ used for forward modeling more closely resemble the field geological structure, we employ the spectral method to generate the conductivity model instead of simply embedding anomalies within a homogeneous half-space underground.
We set the area of interest for conductivity to 200 km $\times$ 100 km (width $\times$ height). We discretize the model into 64 grids along the horizontal direction and 64 grids along the vertical direction, with grid intervals that increase in size. Specifically, we divide the space between 0 km and -1 km into 20 grids at fixed intervals, and discretize the area between -1 km and -20 km and the domain between -20 km and -100 km into 20 and 24 grids, respectively, with intervals that increase logarithmically. To ensure FDM forward modeling fits the boundary conditions of PDEs, we expand the width to 600 km by repeating the conductivity at the margins, extend the depth to -200 km by linearly decaying the conductivity at the bottom, and add an air layer of 600 km $\times$ 200 km. We set the number of grids to 10 for both the expanded area and the air layer. It is worth pointing out that the air layer is only required in mode $xy$.
@@
where $k$ denotes the wavenumber, and $\beta$ represents the scale. Based on common scales of conductivity anomalies found in the earth, we use average conductivity values for five sections of conductivity, with $\beta$ of 3, 4, 5, 6, and 7, respectively. We define the conductivity interval from $10^{-4}$ S/m to 1 S/m, set the conductivity of the air layer to $10^{-9}$ S/m, and the conductivity at the lower boundary to $10^{-2}$ S/m. An example demonstrating the synthetic conductivity model is shown in Fig. 3.
Table 3: Statistics of training and testing datasets. The frequencies and coordinates of TeData-I are same with TrData. The frequencies and coordinates of TeData-II are the downsampled frequencies and coordinates of TrData. The frequencies of TeData-III are different from TrData, and its coordinates are same with TrData.
Dataset TrData TeData-I TeData-II TeData-III
@@
We uniformly arrange 64 sites within the range of $y=-100$ km to $100$ km along the surface at $z=0$ km. We employ the FDM to simulate the electric and magnetic fields. We generate 15,000 pairs ${(\sigma,y,z,f),(\rho_{xy},\rho_{yx},\phi_{xy},\phi_{yx})}$, comprising conductivity model, apparent resistivity, and phase, for training EFKAN at 64 frequencies spaced evenly on a logarithmic scale from 0.049 Hz to 10 Hz, referring to this dataset as TrData. Additionally, we generate three datasets to test the performance of EFKAN: (1) TeData-I. A dataset containing 100 pairs of ${(\sigma,y,z,f),(\rho_{xy},\rho_{yx},\phi_{xy},\phi_{yx})}$ with the same frequencies and spatial coordinates as TrData; (2) TeData-II. This dataset includes 100 resistivity sections with rectangular anomalies embedded in a smooth background, and the electrical and magnetic fields are computed using the same frequencies and spatial coordinates as TrData; (3) TeData-III. This dataset contains 100 resistivity profiles with rectangular anomalies, and the electrical and magnetic fields are simulated using 64 different frequencies (from 0.005 Hz to 12.589 Hz). The statistics of the training and testing datasets are listed in Table 3.
@@
3.2 Prediction by the EFKAN Trained by TrData
In this section, we explore the performance of EFKAN trained by the TrData. For both EFNO and EFKAN, we set the batch size to 50 and the number of training epochs to 200 for both training and testing. We optimize both EFNO and EFKAN using the AdamW optimizer with a learning rate of 0.001, aiming to minimize the relative $\ell_{1}$-norm loss,
@@ 3.2 Prediction by the EFKAN Trained by TrData
where $\hat{\rho}{xy},\hat{\rho}{yx},\hat{\phi}{xy}$, and $\hat{\phi}{yx}$ respectively denote the predicted resistivity and phase for models $xy$ and $yx$, and $N$ indicates the batch size. Additionally, we apply early stopping during training, terminating it when the number of epochs with higher average error on the training samples than on the testing samples exceeds 10. The average loss is calculated as
@@
where $M$ is the number of samples, and $K$ indicates the number of iterations per epoch. The loss curves on TrData and TeData-I are shown in Fig. 4. It can be observed that the loss curves for both EFNO and EFKAN converge rapidly during training. However, the training and testing loss curves of EFNO fluctuate significantly due to early stopping, and the training of EFNO is terminated at the 106-th epoch. Nevertheless, the training and testing loss curves of EFKAN are much smoother and converge to a smaller value than that of EFNO. The quantitative results are listed in Table 4, which proves that EFKAN can achieve higher accuracy than EFNO.
@@
To further explore the capability of EFKAN, we randomly select a resistivity mode from each of the three datasets TeData-I, TeData-II, and TeData-III (Fig. 5). The predicted apparent resistivity and phase for the smooth resistivity (Fig. 5(a)) are shown in Fig. 6, it can be observed that the apparent resistivity and phase predicted by EFNO and EFKAN have high similarities with FDM; however, the error by EFKAN is weaker than that of EFNO. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.0081, 0.0150, 0.0178, and 0.0153, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.0043, 0.0056, 0.0059, and 0.0093, respectively. The 1-D profiles at 0.049 Hz and 10 Hz (Fig. 7) further demonstrate that EFKAN achieves higher accuracy than EFNO.
@@
For the smooth resistivity with rectangular anomalies (Fig. 5(b)) from TeData-II, both EFNO and EFKAN can obtain satisfactory results (Fig. 8). The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.0826, 0.0867, 0.0701, and 0.0852, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.0908, 0.0969, 0.0412, and 0.0758, respectively. These errors are higher than those of TeData-I due to the absence of rectangular anomalies in the training dataset (i.e., TrData). As shown in Fig. 9, although the predictions from EFNO and EFKAN are essentially the same, EFKAN fits the ground truth (i.e., the results by FDM) better, especially in the phase at 10 Hz (Fig. 9(b2) and (b4)).
@@
For the smooth resistivity with rectangular anomalies (Fig. 5(c)) from TeData-III, although the frequencies used for prediction are entirely different from TrData, both EFNO and EFKAN predict the apparent resistivity and phase precisely. The relative $\ell_{1}$-norm errors for $\rho_{xy}$, $\rho_{yx}$, $\phi_{xy}$, and $\phi_{yx}$ by EFNO are 0.0391, 0.0845, 0.0799, and 0.1349, respectively. The relative $\ell_{1}$-norm errors for $\rho_{xy}$, $\rho_{yx}$, $\phi_{xy}$, and $\phi_{yx}$ by EFKAN are 0.0451, 0.0877, 0.0666, and 0.1485, respectively. As shown in Fig. 11, the prediction by EFKAN are closer to the ground truth in some areas, such as the apparent resistivity between 0 and 75 km at 12.589 Hz (Fig. 11(b1) and (b3)).
Table 4: Quantitative results for TrData dataset.
Model TeData-I TeData-II TeData-III
@@
Figure 4: (a) The training loss of EFNO and EFKAN on the TrData dataset; (b) The testing loss of EFNO and EFKAN on the TeData-I dataset.
Figure 5: Examples for demonstrating the effectiveness of EFKAN: (a) The smooth resistivity model randomly sampled from TeData-I; (b) The smooth resistivity model with rectangular anomalies sampled from TeData-II; (c) The smooth resistivity model with rectangular anomalies sampled from TeData-III.
@@
3.3 Prediction by the EFKAN Trained by the TrData with downsampled frequencies and coordinates
@@
In this section, we downsample the frequencies and coordinates of TrData by a factor of 3 to train EFNO and EFKAN, aiming to compare their effectiveness in predicting high-resolution data. Specifically, we use the network trained with apparent resistivity and phase at 22 frequencies and 22 coordinates to predict data at 64 frequencies and 64 coordinates. In Fig. 12, we display the training loss on the downsampled TrData and the testing loss on the TeData-I. The loss curves of EFNO show fluctuations and termination at the 82nd epoch. The training of EFKAN is terminated at the 149-th epoch, and it achieved smaller testing error than EFNO. Table 5 clearly shows that EFKAN achieves smaller average $\ell_{1}$-norm error on TeData-I, TeData-II, and TeData-III, verifying the effectiveness of EFKAN for predicting high-resolution data. Besides, the quantitative results (Table 5) demonstrate that EFKAN outperforms EFNO in terms of prediction accuracy, proving its effectiveness for three different kinds of testing data.
@@
Also, we randomly sample a resistivity model (Fig. 13) from each of the three datasets TeData-I, TeData-II, and TeData-III for illustrating the effectiveness of EFKAN. In Fig. 14, we present the predicted apparent resistivity and phase for the smooth resistivity (Fig. 13(a)). We can observe that the apparent resistivity and phase predicted by EFNO and EFKAN have high precision, and the $\ell_{1}$-norm error of EFKAN is lower than that of EFNO, as corroborated by the color bar. The 1-D profiles at 0.049 Hz and 10 Hz (Fig. 15) further demonstrate that EFKAN fits the ground truth with higher accuracy than EFNO. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.0107, 0.0123, 0.0190, and 0.0210, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.0069, 0.0069, 0.0076, and 0.0110, respectively.
@@
For the smooth resistivity model with rectangular anomalies (Fig. 13(b)) from TeData-II dataset, although the complexity of resistivity has been increased, both EFNO and EFKAN predict the apparent resistivity and phase with high resolution (Fig. 16) and their error is slight. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.0646, 0.1093, 0.0569, and 0.0969, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.0662, 0.0899, 0.0498, and 0.0982, respectively. In the 1-D profiles (Fig. 17), we can further see that the prediction by EFKAN meets the ground truth essentially at 0.049 Hz (Fig. 17(a1$\sim$(a4)) and fits the true apparent resistivity and phase very well at 10 Hz (Fig. 9(b1$\sim$(b4)), outperforming EFNO.
@@
We downsample the original frequencies and coordinates of TeData-III (Fig. 13)(c) to testing the effectiveness of EFKAN for predicting the the apparent resistivity and phase at 64 frequencies and 64 coordinates. As shown in Fig. 18, EFNO and EFKAN present satisfactory predicted results. Additionally, the $\ell_{1}$-norm error of EFKAN is obviously slighter than that of EFNO. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ of EFNO is 0.0397, 0.0856, 0.0802, and 0.1401, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ of EFKAN is 0.0415, 0.0904, 0.0604, and 0.1499, respectively. Fig. 11 shows that the predictions by EFNO and EFKAN are similar at 0.005 Hz, while the predictions by EFKAN are more precise than those of EFNO at 12.589 Hz.
@@
Figure 12: (a) The training loss of EFNO and EFKAN on the downsampled TrData dataset; (b) The testing loss of EFNO and EFKAN on the TeData-I dataset.
Table 5: Quantitative Results for the Down-Sampled TrData.
Model TeData-I TeData-II TeData-III
@@
Figure 13: Examples for demonstrating the effectiveness of EFKAN trained by the downsampled TrData: (a) The smooth resistivity model randomly sampled from TeData-I; (b) The smooth resistivity model with rectangular anomalies sampled from TeData-II; (c) The smooth resistivity model with rectangular anomalies sampled from TeData-III.
@@
3.4 Prediction by the EFKAN Trained by the Small-Scale TrData

### v2

@@ 3 Computational Experiments
In this section, we use synthetic data to evaluate the effectiveness of the proposed method. To make the experiments as close as possible to the field MT forward modeling, we utilize the Gaussian random field (GRF)-based approach to simulate resistivity models for training and testing the EFKAN. We present a series of computational experiments that compare the efficiency and accuracy of EFNO and EFKAN in solving 2-D MT forward problems. Specifically, we focus on using EFKAN to improve the accuracy of solutions for the 2-D Helmholtz equation and explore its applicability to spatial-temporal coordinates and frequencies not included in the training data, as well as its performance on small-scale datasets. All experiments are implemented with the PyTorch platform, and we train and test EFKAN and its competitive approach on a NVIDIA Tesla K80 GPU. It should be noted that the color bars associated with EFKAN errors differ from those associated with EFNO errors in this study. Consequently, when comparing the prediction errors of EFNO and EFKAN, it is essential to take into account the differences between their respective color bars.
@@ 3.1 Data Generation
To make the conductivity $\sigma$ used for forward modeling more closely resemble the geological structure of the field, we employ the spectral method to generate the conductivity model instead of simply embedding anomalies within a homogeneous half-space underground.
We set the conductivity area of interest to 200 km $\times$ 100 km (width $\times$ height). We discretize the model into 64 grids along the horizontal direction and 64 grids along the vertical direction, with grid intervals that increase in size. Specifically, we divide the space between 0 km and -1 km into 20 grids at fixed intervals, and discretize the area between -1 km and -20 km and the domain between -20 km and -100 km into 20 and 24 grids, respectively, with intervals that increase logarithmically. To ensure that FDM forward modeling fits the boundary conditions of PDEs, we expand the width to 600 km by repeating the conductivity at the margins, extend the depth to -200 km by linearly decaying the conductivity at the bottom, and add an air layer of 600 km $\times$ 200 km. We set the number of grids to 10 for both the expanded area and the air layer. It is worth pointing out that the air layer is required only in mode $xy$.
@@
where $k$ denotes the wavenumber and $\beta$ represents the scale. Based on common scales of conductivity anomalies found in the earth, we use average conductivity values for five sections of conductivity, with $\beta$ of 3, 4, 5, 6, and 7, respectively. We define the conductivity interval from $10^{-4}$ S/m to 1 S/m, set the conductivity of the air layer to $10^{-9}$ S/m, and the conductivity at the lower boundary to $10^{-2}$ S/m. An example demonstrating the synthetic conductivity model is shown in Fig. 3.
Table 3: Statistics of training and testing datasets. The frequencies and coordinates of testing dataset A are same with the original training dataset. The number of frequencies and coordinates of testing dataset B are half of those in the original training dataset. The coordinates of C are same with that of the original training dataset, while its frequencies are total different from the original training dataset.
Dataset Original training dataset Testing dataset A Testing dataset B Testing dataset C
@@
We uniformly arrange 64 sites within the range of $y=-100$ km to $100$ km along the surface at $z=0$ km. We employ the FDM to simulate the electric and magnetic fields. We generate 15,000 pairs $\{(\sigma,y,z,f),(\rho_{xy},\rho_{yx},\phi_{xy},\phi_{yx})\}$, comprising the conductivity model, apparent resistivity, and phase, for training EFKAN at 64 frequencies evenly spaced on a logarithmic scale from 0.049 Hz to 10 Hz, referring to this dataset as the original training dataset. Additionally, we generate three datasets to test the performance of EFKAN: (1) Testing dataset A. A has 100 pairs of $\{(\sigma,y,z,f),(\rho_{xy},\rho_{yx},\phi_{xy},\phi_{yx})\}$ with the same frequencies and spatial coordinates as the original training dataset; (2) Testing dataset B. B includes 100 resistivity sections with rectangular anomalies that are unseen in the original training dataset embedded in a smooth background, and the electrical and magnetic fields corresponding to the downsampling frequencies and coordinates of the original training dataset; (3) Testing dataset C. C has 100 resistivity models with rectangular anomalies and the electrical and magnetic fields corresponding to the 64 frequencies (from 0.005 Hz to 12.589 Hz) that are different from those of the original training dataset. The statistics of the training and testing datasets are listed in Table 3. Therefore, we will train a network for each of the following datasets: original training dataset, the dataset with downsampling frequencies and coordinates (i.e., 22 frequencies and 22 coordinates), and the downscaling dataset (i.e., 5000 samples). We will then evaluate each model on datasets A, B, and C.
@@
3.2 Performance on the Original Training Dataset
In this section, we evaluate the EFKAN trained by the original training dataset. For both EFNO and EFKAN training, we set the batch size to 50 and the number of training epochs to 200. We optimize both EFNO and EFKAN using the AdamW optimizer with a learning rate of 0.001 by minimizing the relative $\ell_{1}$-norm loss,
@@ 3.2 Performance on the Original Training Dataset
where $\hat{\rho}{xy},\hat{\rho}{yx},\hat{\phi}{xy}$, and $\hat{\phi}{yx}$ denote the predicted resistivity and phase for models $xy$ and $yx$, respectively, and $N$ indicates the batch size. Additionally, we apply early stopping for network training, terminating it when the number of epochs with a higher average error than the testing error exceeds 10. The average loss is expressed as
@@
where $M$ is the number of samples and $K$ indicates the number of iteration per epoch. The loss curves on the original training dataset and testing dataset A are shown in Fig. 4. It can be observed that the loss curves for both EFNO and EFKAN rapidly converge. However, the training and testing loss curves of EFNO fluctuate significantly, and the training of EFNO is terminated at the 106-th epoch. The training and testing loss curves of EFKAN are much smoother and converge to a smaller value than those of EFNO. The quantitative results are listed in Table 4, which proves that EFKAN can achieve higher precision than EFNO.
@@
We further randomly select a resistivity mode from each of testing dataset A, B, and C (Fig. 5). The predicted apparent resistivity and phase for the smooth resistivity (Fig. 5(a)) are shown in Fig. 6, it can be observed that both apparent resistivity and phase predicted by EFNO and EFKAN have high similarities with FDM, and the error by EFKAN is smaller than that of EFNO. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.0081, 0.0150, 0.0178, and 0.0153, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.0043, 0.0056, 0.0059, and 0.0093, respectively. The 1-D profiles at 0.049 Hz and 10 Hz (Fig. 7) further demonstrate that EFKAN has high precision for solving PDEs.
@@
For the smooth resistivity with rectangular anomalies (Fig. 5(b)) from testing dataset B, both EFNO and EFKAN can obtain satisfactory results (Fig. 8). The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.0826, 0.0867, 0.0701, and 0.0852, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.0908, 0.0969, 0.0412, and 0.0758, respectively. These errors are higher than those of testing dataset A due to the original training dataset having no rectangular anomalies. As shown in Fig. 9, although the predictions from EFNO and EFKAN are essentially the same, EFKAN fits the ground truth (i.e. the results by FDM) better, especially in the phase at 10 Hz (Fig. 9(b2) and (b4)).
@@
For the smooth resistivity with rectangular anomalies (Fig. 5(c)) from testing dataset C, although the frequencies used for prediction are entirely different from the original training dataset, both EFNO and EFKAN predict the apparent resistivity and phase precisely. The relative $\ell_{1}$-norm errors for $\rho_{xy}$, $\rho_{yx}$, $\phi_{xy}$, and $\phi_{yx}$ by EFNO are 0.0391, 0.0845, 0.0799, and 0.1349, respectively. The relative $\ell_{1}$-norm errors for $\rho_{xy}$, $\rho_{yx}$, $\phi_{xy}$, and $\phi_{yx}$ by EFKAN are 0.0451, 0.0877, 0.0666, and 0.1485, respectively. As shown in Fig. 11, the predictions by EFKAN have high precision with the ground truth in certain areas, such as the apparent resistivity between 0 and 75 km at 12.589 Hz (Fig. 11(b1) and (b3)).
Table 4: Quantitative results for the original training dataset dataset.
Model Testing dataset A Testing dataset B Testing dataset C
@@
Figure 4: (a) The training loss of EFNO and EFKAN on the the original training dataset dataset; (b) The testing loss of EFNO and EFKAN on testing dataset A.
Figure 5: Examples for demonstrating the effectiveness of EFKAN: (a) The smooth resistivity model randomly sampled from testing dataset A; (b) The smooth resistivity model with rectangular anomalies sampled from testing dataset B; (c) The smooth resistivity model with rectangular anomalies sampled from testing dataset C.
@@
3.3 Performance on the Downsampling Training Dataset
@@
In this section, we downsample both the frequencies and coordinates of the original training dataset by a factor of 3 to train EFNO and EFKAN, aiming to compare their effectiveness in predicting high-resolution data. Specifically, we use the network trained with apparent resistivity and phase at 22 frequencies and 22 coordinates to predict data at 64 frequencies and 64 coordinates. In Fig. 12, we display the training loss in the downsampling original training dataset and the testing loss on testing dataset A. The loss curves of EFNO show fluctuations and stopped at the 82nd epoch. The EFKAN training terminated at the 149-th epoch, and it achieved a smaller error than EFNO. Table 5 clearly shows that EFKAN achieves a smaller average $\ell_{1}$-norm error on testing dataset A, B, and C, which verifies the effectiveness of EFKAN in predicting high-resolution data. In addition, the quantitative results (Table 5) demonstrate that EFKAN outperforms EFNO in terms of prediction accuracy, demonstrating its effectiveness for three different kinds of testing data.
@@
In addition, we randomly sample a resistivity model (Fig. 13) from each of testing dataset A, B, and C to show the effectiveness of EFKAN. In Fig. 14, we present the predicted apparent resistivity and phase for smooth resistivity (Fig. 13(a)). We can observe that the apparent resistivity and phase predicted by EFNO and EFKAN have high precision, and the $\ell_{1}$-norm error of EFKAN is lower than that of EFNO, as corroborated by the color bar. The 1-D profiles at 0.049 Hz and 10 Hz (Fig. 15) further demonstrate that EFKAN can fit the ground truth with higher accuracy than EFNO. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.0107, 0.0123, 0.0190, and 0.0210, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.0069, 0.0069, 0.0076, and 0.0110, respectively.
@@
For the smooth resistivity model with rectangular anomalies (Fig. 13(b)) from testing dataset B, although the complexity of the resistivity has increased, both EFNO and EFKAN predict the apparent resistivity and phase with high resolution (Fig. 16). The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.0646, 0.1093, 0.0569, and 0.0969, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.0662, 0.0899, 0.0498, and 0.0982, respectively. In the 1-D profiles (Fig. 17), we can further observe that the prediction by EFKAN meets the ground truth essentially at 0.049 Hz (Fig. 17(a1$\sim$(a4)) and fits the true apparent resistivity and phase very well at 10 Hz (Fig. 17(b1$\sim$(b4)), outperforming EFNO.
@@
For the prediction of the frequencies different from those of the training dataset, Fig. 18 shows the predictions of the smooth resistivity model with rectangular anomalies (Fig. 13(c)) by both EFNO and EFKAN present satisfactory. The $\ell_{1}$-norm error of EFKAN is lower than that of EFNO. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ of EFNO is 0.0397, 0.0856, 0.0802, and 0.1401, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ of EFKAN is 0.0415, 0.0904, 0.0604, and 0.1499, respectively. Fig. 19 shows that the solutions of EFNO and EFKAN are similar at 0.005 Hz, while the predictions of EFKAN are more precise than those of EFNO at 12.589 Hz.
@@
Figure 12: (a) The training loss of EFNO and EFKAN on the downsampling original training dataset dataset; (b) The testing loss of EFNO and EFKAN on testing dataset A.
Table 5: Quantitative Results for the downsampling original training dataset.
Model Testing dataset A Testing dataset B Testing dataset C
@@
Figure 13: Examples for demonstrating the effectiveness of EFKAN trained by the downsampling original training dataset: (a) The smooth resistivity model randomly sampled from testing dataset A; (b) The smooth resistivity model with rectangular anomalies sampled from testing dataset B; (c) The smooth resistivity model with rectangular anomalies sampled from testing dataset C.
@@
3.4 Performance on the Downscaling Training Dataset